# CAD-Editor: A Locate-then-Infill Framework with Automated Training Data Synthesis for Text-Based CAD Editing

### v1

@@ Abstract
Computer Aided Design (CAD) is indispensable across various industries. Text-based CAD editing, which automates the modification of CAD models based on textual instructions, holds great potential but remains underexplored. Existing methods primarily focus on design variation generation or text-based CAD generation, either lacking support for text-based control or neglecting existing CAD models as constraints. We introduce CAD-Editor, the first framework for text-based CAD editing. To address the challenge of demanding triplet data with accurate correspondence for training, we propose an automated data synthesis pipeline. This pipeline utilizes design variation models to generate pairs of original and edited CAD models and employs Large Vision-Language Models (LVLMs) to summarize their differences into editing instructions. To tackle the composite nature of text-based CAD editing, we propose a locate-then-infill framework that decomposes the task into two focused sub-tasks: locating regions requiring modification and infilling these regions with appropriate edits. Large Language Models (LLMs) serve as the backbone for both sub-tasks, leveraging their capabilities in natural language understanding and CAD knowledge. Experiments show that CAD-Editor achieves superior performance both quantitatively and qualitatively.
@@ 1 Introduction
CAD model creation is an iterative process, where an initial draft undergoes multiple modifications until it aligns with user requirements. Natural language plays a crucial role throughout this process, serving as a key medium of communication. For non-experts, it offers the most intuitive way to express their needs, while for professionals, it enables fast, detailed, and precise instructions. Consequently, a system capable of automatically editing CAD models based on textual instructions – known as text-based CAD editing (Figure 1) – has the potential to revolutionize the entire CAD design workflow. Such a system could significantly accelerate CAD model development and empower a broader range of individuals, especially those with limited design expertise, to create CAD models more effectively.
@@
While important, text-based CAD editing receives limited attention. Some studies explore design variation generation, where new CAD models are generated by randomly altering components of an existing model (Wu et al., 2021; Xu et al., 2022, 2023; Zhang et al., 2024b). However, these approaches lack support for text-based control over the appearance of the generated CAD models, limiting their practical usability. Another line of research makes initial attempts at text-based CAD generation, focusing on generating new CAD models directly from textual descriptions (Khan et al., 2024b). Nonetheless, these methods do not incorporate an existing CAD model as input, which prevents them from leveraging the original design’s context and constraints.
@@
Text-based CAD editing presents several distinct challenges. First, training for this task requires triplet data with accurate correspondence among an original CAD model, an editing instruction and an edited CAD model. However, such data does not naturally exist, and manually collection is both costly and difficult to scale. Second, text-based CAD editing is inherently a composite problem. It demands a comprehensive understanding of the textual description, the ability to locate the corresponding parts within the intricate structure of the CAD’s SE operations, and the capability to generate concrete modifications to these SE operations.
@@
In this work, we introduce CAD-Editor, the first framework for text-based CAD editing. We frame the task as a sequence-to-sequence (seq2seq) generation problem, where the input combines an editing instruction and the sequence representation of the original CAD model, and the output is the sequence of the edited CAD model (Figure 2). To address the need for triplet data with accurate correspondence, we propose an automated data synthesis pipeline that leverages existing design variation models and Large Vision-Language Models (LVLMs) (Figure 3). Starting from an existing CAD model, design variation models generate edited CAD models by randomly altering parts while keeping others unchanged, producing pairs of original and edited CAD models. LVLMs then summarize the differences between these CAD models into editing instructions, resulting in triplets with strong correspondence. To tackle the composite nature of text-based CAD editing, we decompose the task into specialized sub-tasks: locating and infilling, each addressing a specific aspect of the editing process (Figure 4). For both sub-tasks, Large Language Models (LLMs) serve as the backbone, leveraging their natural language understanding and CAD design knowledge (Makatura et al., 2023). In the locating stage, LLMs identify regions requiring modification by generating a masked CAD sequence, where special tokens <mask> indicate the regions to be modified. In the infilling stage, LLMs generate appropriate edits for these masked regions, using the masked CAD sequence from the locating stage as context.
@@
  • We develop a locate-then-infill framework that decomposes the task into focused sub-tasks, effectively handling the composite nature of text-based CAD editing.
@@ 2 Related Works
CAD Generation. Parametric CAD, defined by its sketch-and-extrude operations, is central to mechanical design due to its ability to retain modeling history, which facilitates both editing and manufacturing. Recent large-scale CAD datasets (Wu et al., 2021) have fueled the development of generative models. Wu et al. (2021) explores unconditional generation, where a random latent vector is used as input to generate CAD models. Xu et al. (2022, 2023); Zhang et al. (2024b) focus on design variation generation, which randomly modifies parts of an existing CAD model. Khan et al. (2024a) studies text-based CAD generation, which transforms textual descriptions into CAD models. Our work differs in two key aspects. First, we target a distinct task, unlike prior work, which either lacks text-based control (Wu et al., 2021; Xu et al., 2022, 2023; Zhang et al., 2024b) or disregards existing CAD models as constraints (Khan et al., 2024a). Second, we introduce a novel locate-then-infill framework based on LLMs to handle the composite nature of text-based CAD editing. Previous approaches either rely on a VAE-based framework (Wu et al., 2021; Xu et al., 2022, 2023) or use LLMs without accounting for task-specific needs (Zhang et al., 2024b; Khan et al., 2024a).
Large Language Models (LLMs). Recently, LLMs like GPT (Achiam et al., 2023; OpenAI, 2023) and LLaMA (Touvron et al., 2023) have distinguished themselves from smaller models through remarkable emergent capabilities. Beyond excelling in natural language processing, LLMs have transforms generative tasks in other domains, e.g., motion generation (Zhang et al., 2024a) and material generation (Gruver et al., 2024). These advancements inspire us to adopt LLMs as the backbone for sub-tasks in our locate-then-infill framework. Moreover, LLMs and Large Vision-Language Models (LVLMs) are increasingly utilized for data synthesis to enhance training (Xu et al., 2024; Yu et al., 2024). Specifically, Khan et al. (2024b) leverage LLMs/LVLMs to synthesize data for text-based CAD generation. However, our task of text-based CAD editing presents distinct challenges. First, unlike Khan et al. (2024a), who generate two-tuple data (a text prompt and a CAD model), our task involves creating triplet data: an editing instruction, an original CAD model, and an edited CAD model. We address this by combining design variation models with LLMs/LVLMs. Second, while Khan et al. (2024a) focus on captioning single CAD models, our task requires summarizing differences between two CAD models. We handle this by introducing a stepwise captioning strategy.
@@ 5.3 Training and Inference
Inference. The locating and the infilling stage operates sequentially. The locating stage generates $\mathcal{C}_{\text{mask}}$ using $\mathcal{C}_{\text{orig}}$ and $\mathcal{I}$ as input. The infilling stage generates $\mathcal{C}_{\text{edit}}$ by using $\mathcal{C}_{\text{mask}}$ from the locating stage along with $\mathcal{C}_{\text{orig}}$ and $\mathcal{I}$ as input.
@@ 6 Experiments
Figure 5: Qualitative results from CAD-Editor, GPT-4o-Basic and GPT-4o-IC . The text below shows the editing instruction.
@@ 6.1 Experimental Setup
Datasets. We use the DeepCAD dataset (Wu et al., 2021) which contains 178k CAD models. We split it into 90% training, 5% validation, and 5% testing segments. We follow the same strategy in existing work (Xu et al., 2022, 2023) to remove duplication. Besides, we also exclude the long-tail data with more than 3 sketch-extrusion pairs. For the synthetic dataset used for training, we generate 120k examples using the method introduced in Section 4. For the test set, we randomly sample 2k examples from the original test segment, generate the initial version following Section 4, and manually examine them to ensure the correctness. To compare the performance of different methods, we generate 5 outputs for each example in the test set, yielding 10k CAD models for evaluation.
@@
Baselines. We compare our results with three types of baselines: 1) design variation generation models, including SkexGen (Xu et al., 2022), Hnc-CAD (Xu et al., 2023) and FlexCAD (Zhang et al., 2024b); 2) text-based CAD generation models such as Text2CAD (Khan et al., 2024a) and 3) foundation models that are not specifically designed for CAD generation but have acquired some CAD knowledge during pre-training. For the third category, we select one of the most powerful foundation models, GPT-4o, as our baseline. We use two prompting strategies: 1) GPT-4o-Basic, which provides only an explanation of CAD operation sequences; and 2) GPT-4o-IC: which includes three in-context (IC) examples retrieved based on cosine similarity between editing instructions, in addition to the basic explanation. Notably, existing CAD design variation models and text-based CAD generation models do not support text-based CAD editing. We include them as baselines to show that our model can generate CAD models with comparable or superior quality while addressing a more complex task.
Table 1: Quantitative evaluations. SkexGen, Hnc-CAD, FlexCAD and Text2CAD do not support text-based editing, so only their generation quality is compared. JSD and D-CLIP values are scaled by $10^{2}$. $\uparrow$: the higher the better, $\downarrow$: the lower the better.
Method JSD $\downarrow$ D-CLIP $\uparrow$ VR $\uparrow$ H-Eval $\uparrow$
SkexGen 1.94 - 74.3 -
Hnc-CAD 1.77 - 77.4 -
FlexCAD 1.72 - 82.1 -
Text2CAD 2.39 - 84.8 -
GPT-4o-Basic 1.10 - 1.08 63.2 7.22
GPT-4o-IC 0.70 - 0.11 84.5 15.6
CAD-Editor 0.65 0.27 95.6 43.2
@@
Metrics. As this work is the first to address text-based CAD editing, we propose evaluating the task from three key aspects. 1) Validity. The generated CAD sequence must be valid, meaning it can be successfully parsed and rendered into a visual object. We denote this as Valid Ratio (VR). 2) Realism. The generated CAD models should be realistic and similar to ground-truth CAD models. To measure this, we adopt the Jensen-Shannon Divergence (JSD) from prior work (Wu et al., 2021; Xu et al., 2022, 2023). JSD quantifies the similarity between two probability distributions, evaluating how often the ground-truth point clouds occupy similar positions as the generated point clouds. 3) Edit Consistency. The modifications in the generated CAD model should align with the given editing instructions. To assess this, we adapt Directional CLIP Score (D-CLIP) from the image generation domain (Gal et al., 2022) as the metric. This metric evaluates how well the change between the image for the edited CAD model and the image for the original CAD model aligns with the editing instruction:
@@
$\Delta T=E_{T}\left(t_{\text{edit}}\right)-E_{T}\left(t_{\text{orig}}\right),\Delta I=E_{I}\left(i_{\text{edit}}\right)-E_{I}\left(i_{\text{orig}}\right).$
$E_{I}$ and $E_{T}$ are CLIP’s image and text encoders, $t_{\text{orig}}$ is a neutral text (e.g.,“This is a 3D shape.”), $t_{\text{edit}}$ is the concatenation of $t_{\text{orig}}$ and the textual instruction. $i_{\text{edit }}$ and $i_{\text{orig }}$ are the images for the edited and original CAD model.
@@
6.2 Main Results
Quantitative Evaluation. Table 1 reports the average scores across 3 runs. Notably, CAD-Editor achieves a high Valid Ratio of 95.6%, significantly surpassing other methods and indicating a greater proportion of valid and high-quality CAD generations. In terms of D-CLIP, which measures alignment with editing instructions, CAD-Editor scores 0.27, a substantial improvement over GPT-4o-Basic and GPT-4o-IC. This underscores CAD-Editor’s effectiveness in adhering to user instructions. Additionally, CAD-Editor outperforms all baselines on the point cloud evaluation metric JSD, demonstrating good generation quality. Overall, the results indicate that CAD-Editor not only enable precise text-based CAD editing, which better alignment with user instructions, but also achieve higher validity and better quality of the generated CAD designs.
Besides, we present more results from CAD-Editor. Figure 6 shows how CAD-Editor handles a variety of editing instructions. Figure 7 illustrates that, given the same original CAD model, CAD-Editor applies different modifications based on the provided editing instructions. Figure 8 shows that when the editing instruction is vague, CAD-Editor generates multiple CAD models, all aligning with the user’s intent. Figure 9 highlights the iterative editing capability of CAD-Editor, allowing users to refine a CAD model through successive instructions until a satisfactory result is achieved.

### v2

@@ Abstract
Computer Aided Design (CAD) is indispensable across various industries. Text-based CAD editing, which automates the modification of CAD models based on textual instructions, holds great potential but remains underexplored. Existing methods primarily focus on design variation generation or text-based CAD generation, either lacking support for text-based control or neglecting existing CAD models as constraints. We introduce CAD-Editor, the first framework for text-based CAD editing. To address the challenge of demanding triplet data with accurate correspondence for training, we propose an automated data synthesis pipeline. This pipeline utilizes design variation models to generate pairs of original and edited CAD models and employs Large Vision-Language Models (LVLMs) to summarize their differences into editing instructions. To tackle the composite nature of text-based CAD editing, we propose a locate-then-infill framework that decomposes the task into two focused sub-tasks: locating regions requiring modification and infilling these regions with appropriate edits. Large Language Models (LLMs) serve as the backbone for both sub-tasks, leveraging their capabilities in natural language understanding and CAD knowledge. Experiments show that CAD-Editor achieves superior performance both quantitatively and qualitatively. The code is available at https://github.com/microsoft/CAD-Editor.
@@ 1 Introduction
CAD model creation is an iterative process, where an initial draft undergoes multiple modifications until it aligns with user requirements. Natural language plays a crucial role throughout this process, serving as a key medium of communication. For non-experts, it offers the most intuitive way to express their needs, while for professionals, it enables fast, detailed, and precise instructions. Consequently, a system capable of automatically editing CAD models based on textual instructions – known as text-based CAD editing (Figure 1) – has the potential to revolutionize the entire CAD design workflow. Such a system could significantly accelerate the development of CAD models and empower a broader range of individuals, especially those with limited design expertise, to create CAD models more effectively.
@@
While important, text-based CAD editing receives limited attention. Some studies explore design variation generation, where new CAD models are generated by randomly altering components of an existing model (Wu et al., 2021; Xu et al., 2022, 2023; Zhang et al., 2024b). However, these approaches lack support for text-based control over the appearance of the generated CAD models, limiting their practical usability. Another line of research makes initial attempts at text-based CAD generation, focusing on generating new CAD models directly from textual descriptions (Khan et al., 2024b; Li et al., 2024). Nonetheless, these methods do not incorporate an existing CAD model as input, which prevents them from leveraging the original design’s context and constraints.
@@
Text-based CAD editing presents several distinct challenges. First, training for this task requires triplet data with accurate correspondence among an original CAD model, an editing instruction, and an edited CAD model. However, such data does not naturally exist, and manually collection is both costly and difficult to scale. Second, text-based CAD editing is inherently a composite problem. It demands a comprehensive understanding of diverse textual instructions and geometric concepts, the ability to locate the corresponding parts within the intricate structure of the CAD’s SE operations, and the capability to generate concrete modifications to these SE operations.
@@
In this work, we introduce CAD-Editor, the first framework for text-based CAD editing. We frame the task as a sequence-to-sequence (seq2seq) generation problem, where the input combines an editing instruction and the sequence representation of the original CAD model, and the output is the sequence of the edited CAD model (Figure 2). To address the need for triplet data with accurate correspondence, we propose an automated data synthesis pipeline that leverages existing design variation models and Large Vision-Language Models (LVLMs) (Figure 3). Starting from an existing CAD model, design variation models generate edited CAD models by randomly altering parts while keeping others unchanged, producing pairs of original and edited CAD models. LVLMs then summarize the differences between these CAD models into editing instructions, resulting in triplets with strong correspondence. To tackle the composite nature of text-based CAD editing, we decompose the task into specialized sub-tasks: locating and infilling, each addressing a specific aspect of the editing process (Figure 4). For both sub-tasks, Large Language Models (LLMs) serve as the backbone, leveraging their strong natural language understanding capabilities and basic CAD-related knowledge, including SE operations and geometric concepts (Makatura et al., 2023). In the locating stage, LLMs identify regions requiring modification by generating a masked CAD sequence, where special tokens <mask> indicate the regions to be modified. In the infilling stage, LLMs generate appropriate edits for these masked regions, using the masked CAD sequence from the locating stage as context.
@@
  • We develop a locate-then-infill framework that decomposes the task into focused sub-tasks, and leverage LLMs as the backbone, effectively handling the composite nature of text-based CAD editing.
@@ 2 Related Works
CAD Generation. Parametric CAD (Shahin, 2008; Camba et al., 2016), defined by its sketch-and-extrude operations, is central to mechanical design due to its ability to retain modeling history, which facilitates both editing and manufacturing. Recent large-scale CAD datasets (Wu et al., 2021) have fueled the development of generative models. Wu et al. (2021) explored unconditional generation, where a random latent vector is used as input to generate CAD models. Xu et al. (2022, 2023); Zhang et al. (2024b) focused on design variation generation, which randomly modifies specific part of an existing CAD model. Recently, Khan et al. (2024a); Li et al. (2024) studied text-based CAD generation, which transforms textual descriptions into CAD models. Our work differs in two key aspects. First, we target a distinct task, unlike prior work, which either lacks text-based control (Wu et al., 2021; Xu et al., 2022, 2023; Zhang et al., 2024b) or disregards existing CAD models as constraints (Khan et al., 2024a; Li et al., 2024). Second, we introduce a novel locate-then-infill framework based on LLMs to handle the composite nature of text-based CAD editing. Previous approaches either rely on VAE-based (Wu et al., 2021; Xu et al., 2022, 2023) or transformer-based (Khan et al., 2024a; Li et al., 2024) architectures, or apply LLMs without accounting for task-specific needs (Zhang et al., 2024b).
Large Language Models (LLMs). Recently, LLMs like GPT (Achiam et al., 2023; OpenAI, 2023) and LLaMA (Touvron et al., 2023) have distinguished themselves from smaller models, particularly through advanced prompting (Brown et al., 2020; Wei et al., 2022b; Kojima et al., 2022) or fine-tuning methods (Wei et al., 2022a). Beyond excelling in natural language processing (Yue et al., 2024; Zhan et al., 2025; Zhao et al., 2025), LLMs have transform generative tasks in other domains, e.g. motion generation (Zhang et al., 2024a) and material generation (Gruver et al., 2024). These advancements inspire us to adopt LLMs as the backbone for subtasks in our locate-then-infill framework. Moreover, LLMs and LVLMs are increasingly utilized for data synthesis to enhance training (Xu et al., 2024; Yu et al., 2024). Specifically, Khan et al. (2024b) leverage LLMs/LVLMs to synthesize data for text-based CAD generation. However, our task of text-based CAD editing presents distinct challenges. First, unlike Khan et al. (2024a), who generates two-tuple data (a text prompt and a CAD model), our task involves creating triplet data: an editing instruction, an original CAD model, and an edited CAD model. We address this by combining design variation models with LLMs/LVLMs. Second, while Khan et al. (2024a) focus on captioning single CAD models, our task requires summarizing differences between two CAD models. We handle this by introducing a stepwise captioning strategy.
@@ 5.3 Training and Inference
Inference. During inference, the locating and the infilling stage operates sequentially. The locating stage generates $\mathcal{C}_{\text{mask}}$ using $\mathcal{C}_{\text{orig}}$ and $\mathcal{I}$ as input. The infilling stage generates $\mathcal{C}_{\text{edit}}$ by using $\mathcal{C}_{\text{mask}}$ from the locating stage along with $\mathcal{C}_{\text{orig}}$ and $\mathcal{I}$ as input.
@@ 6 Experiments
Figure 5: Qualitative results from CAD-Editor, GPT-4o-Basic and GPT-4o-IC .
@@ 6.1 Experimental Setup
Datasets. We use the DeepCAD dataset (Wu et al., 2021) which contains 178k CAD models. We split it into 90% training, 5% validation, and 5% testing segments. We follow the same strategy in existing work (Xu et al., 2022, 2023) to remove duplication. For the synthetic dataset used for training, we generate 120k examples using the method introduced in Section 4. For the test set, we randomly sample 2k examples from the original test segment, generate the initial version following Section 4, and manually examine them to ensure the correctness. To compare the performance of different methods, we generate 5 outputs for each example in the test set, yielding 10k CAD models for evaluation.
@@
Baselines. We compare our results with three types of baselines: 1) design variation generation models, including SkexGen (Xu et al., 2022), Hnc-CAD (Xu et al., 2023) and FlexCAD (Zhang et al., 2024b); 2) text-based CAD generation models such as Text2CAD (Khan et al., 2024a) and 3) foundation models that are not specifically designed for CAD generation but have acquired some CAD knowledge during pre-training. For the third category, we select one of the most powerful foundation models, GPT-4o, as our baseline. We use two prompting strategies: 1) GPT-4o-Basic, which provides only an explanation of CAD operation sequences; and 2) GPT-4o-IC: which includes three in-context (IC) examples retrieved based on cosine similarity between editing instructions, in addition to the basic explanation. Notably, existing CAD design variation models and text-based CAD generation models do not support text-based CAD editing. We include them as baselines to show that our model can generate CAD models with comparable or superior validity and quality while addressing a more complex task.
@@
Figure 7: CAD-Editor can deal with parametric instructions, including explicit numeric expressions and implicit parametric cues.
6.2 Metrics
As this work is the first to address text-based CAD editing, we propose evaluating the task from three key aspects.
(1) Validity. The generated CAD sequence must be valid, that is, it can be successfully parsed and rendered into a 3D visual object. We denote this as Valid Ratio (VR).
(2) Realism. The generated CAD models should be realistic and similar to ground-truth CAD models. To measure this, we adopt the Jensen-Shannon Divergence (JSD) from prior work (Wu et al., 2021; Xu et al., 2022, 2023). JSD quantifies the similarity between two probability distributions, evaluating how often the ground-truth point clouds occupy similar positions as the generated point clouds.
(3) Edit Consistency. The generated edits should faithfully reflect the provided textual instructions. We assess consistency at both the point cloud and image levels:
$\bullet$ Point Cloud Level. We use the Chamfer Distance (CD) as a reconstruction metric as used in (Khan et al., 2024b) to evaluate geometric alignment between the edited and ground-truth CAD model pairs. While CD captures shape similarity, it may not fully reflect semantic consistency, as the goal of editing is not exact replication but meaningful transformation aligned with the instruction.
$\bullet$ Image Level. We adapt Directional CLIP Score (D-CLIP) from the image editing domain (Gal et al., 2022; Brooks et al., 2023) as the metric. Building upon CLIP score (Radford et al., 2021; Sohn et al., 2023), D-CLIP evaluates how well the change between the image for the edited CAD model and the image for the original CAD model aligns with the editing instruction:
@@ 6.2 Metrics
$\Delta T=E_{T}\left(t_{\text{edit}}\right)-E_{T}\left(t_{\text{orig}}\right),\Delta I=E_{I}\left(i_{\text{edit}}\right)-E_{I}\left(i_{\text{orig}}\right),$
where $E_{I}$ and $E_{T}$ are CLIP’s image and text encoders, respectively. $t_{\text{orig}}$ is a neutral text (e.g.,“This is a 3D shape.”), $t_{\text{edit}}$ is the concatenation of $t_{\text{orig}}$ and the textual instruction. $i_{\text{edit }}$ and $i_{\text{orig }}$ are the images for the edited and original CAD model, respectively.
Table 1: Quantitative evaluations. SkexGen, Hnc-CAD, FlexCAD and Text2CAD do not support text-based editing, so only their generation quality is compared. JSD, CD, and D-CLIP values are scaled by $10^{2}$. $\uparrow$: the higher the better, $\downarrow$: the lower the better.
Method VR $\uparrow$ JSD $\downarrow$ CD $\downarrow$ D-CLIP $\uparrow$ H-Eval $\uparrow$
SkexGen 74.3 1.94 - - -
Hnc-CAD 77.4 1.77 - - -
FlexCAD 82.1 1.72 - - -
Text2CAD 84.8 2.39 1.91 - -
GPT-4o-Basic 63.2 1.10 2.30 - 1.08 7.22
GPT-4o-IC 84.5 0.70 1.55 - 0.11 15.6
CAD-Editor 95.6 0.65 1.18 0.11 43.2
6.3 Main Results
Quantitative Evaluation. Table 1 reports the average scores across 3 runs. Notably, CAD-Editor achieves a high Valid Ratio of 95.6%, significantly surpassing other methods and indicating a greater proportion of valid and high-quality CAD generations. In terms of CD and D-CLIP, which measure alignment with editing instructions, CAD-Editor achieves scores of 1.18 and 0.11 respectively, representing substantial improvements over both GPT-4o-Basic and GPT-4o-IC. These results underscore CAD-Editor’s effectiveness in adhering to user instructions. Additionally, CAD-Editor outperforms all baselines on the point cloud evaluation metric JSD, demonstrating good generation quality. Overall, the results indicate that CAD-Editor not only enable precise text-based CAD editing, which better alignment with user instructions, but also achieve higher validity and better quality of the generated CAD designs.
Besides, we present more results from CAD-Editor. Figure 6 shows how CAD-Editor handles a variety of editing instructions and different CAD models. Figure 7 shows that CAD-Editor has the ability to interpret parameterized instructions, including explicit numeric expressions such as “add a 44-unit diameter circular hole” and implicit parametric cues like “reduce the cylinder’s height by half”. Figure 8 illustrates that, given the same original CAD model, CAD-Editor applies different modifications based on the provided editing instructions. Figure 9 shows that when the editing instruction is vague, CAD-Editor generates multiple CAD models, all aligning with the user’s intent. Figure 10 highlights the iterative editing capability of CAD-Editor, allowing users to refine a CAD model through successive instructions until a satisfactory result is achieved.
@@ 6.3 Main Results
Method VR $\uparrow$ JSD $\downarrow$ CD $\downarrow$ D-CLIP $\uparrow$
CAD-Editor-mini w/ Basic 88.8 0.78 1.25 - 0.12
CAD-Editor-mini w/ Step 90.1 0.70 1.22 - 0.07
CAD-Editor w/ Direct 86.1 0.77 1.26 - 0.19
CAD-Editor w/ L-I 96.5 0.67 1.13 0.03
CAD-Editor w/ L-I & HS 95.6 0.65 1.18 0.11